OpenStudy (anonymous):

A survey of 100 randomly selected students gave the following information: 45 are taking math 41 are taking English 40 are taking history 15 are taking math and English 18 are taking math and history 17 are taking English and history 7 are taking all three How many students are only taking mathematics?

OpenStudy (geometry_hater):

ONLY so im going with 45

OpenStudy (amistre64):

nope, this can be visually seen with venss, but the mathing still works the same; start with the "all three" and peel your way out

OpenStudy (amistre64):

7 are taking all three; MnEnH = 7 15 are taking math and English; MnE = 15-7 = 8 18 are taking math and history; MnH = 18-7 = 9 17 are taking English and history; EnH = 17-7=10  45 are taking math; but we already have 7+8+9 accounted for 45 - 7 - 8 - 9 = those only taking math